# Zeal: Rethinking Large-Scale Resource Allocation with “Decouple and Decompose”

### v1

@@ 7.1.2 Cluster Scheduling
Max-min allocation.  Figure 5 compares Zeal against other approaches with the objective of max-min allocation. Zeal achieves a near-optimal max-min allocation of 0.97 within 5 seconds and 0.98 within 10 seconds, while Zeal reaches 0.98 within 2 seconds. Due to their sequential nature, Exact Sol and Gandiva encounter a slow computation time of 86 and 22 seconds. The greedy heuristic Gandiva even experiences a 0.56 low max-min allocation. POP-4 deliverers a comparable time-accuracy tradeoff as Zeal but is 3.4$\times$ slower than Zeal, which is simulated in the same way as POP did for a fair comparison. A faster POP variant, POP-16, exhibits a lower max-min allocation of 0.94, as it becomes more difficult for jobs to assess the ideal GPU/CPU among the limited selection in smaller subproblems.
@@
Proportional fairness.  Figure 6 compares Zeal against other approaches in maximizing proportional fairness. As the optimization problem becomes more difficult with the convex logarithm function in the objective, both Zeal and Zeal achieve a greater speedup than POP by thoroughly breaking down the problem into smaller and easier pieces. Zeal and Zeal reach a proportional fairness score of over 1,000 in 300 seconds and 100 seconds, respectively. In contrast, POP-4 takes 1,825 seconds to reach the same score, while POP-16 and POP-64 remain below 900 despite their shorter computation times. Exact Sol, though accurate, is prohibitively slow, requiring 12,432 seconds.
@@ 7.1.3 Load Balancing
Minimize shard movements.  Figure 8 compares Zeal with state-of-the-art approaches in minimizing shard movements. Zeal achieves an average of 24.1 shard movements in 15 seconds, slightly better than POP-4’s 25.4 shard movements within 16.9 seconds. In faster scenarios, Zeal achieves 35.1 shard movements in 2 seconds, while POP-4 achieves slightly fewer movements of 32.2 in 1.7 seconds. In a more fair comparison with POP, Zeal consistently costs 15–18% fewer shard movements than POP given a similar time limit. Exact Sol achieves the minimum shard movements of 13.2 but is significantly slower, taking an average of 1,572 seconds due to the NP-hardness of solving the mixed-integer linear programming problem inherent to load balancing. On the other hand, the greedy heuristic represents the opposite end of the time-accuracy trade-off, completing in just 25 milliseconds but requiring 50 shard movements.
@@ 7.2 Robustness
We evaluate the robustness of Zeal and other baselines against changes in problem granularity, temporal fluctuations, spatial redistribution, and link failures. In all these evaluations, Zeal is run for 30 seconds. To mitigate the impact of the changing resources and demands, we normalize the satisfied demand by the optimal satisfied demand from Exact Sol.
@@
As the mean edge betweenness centrality decreases, all approaches experience a decline in normalized satisfied demand, ranging from 0.8–2.1% for Zeal, 2.8–4.2% for Teal, 2.5–5.6% for Pinning, and 1–12.3% for POP. However, the maximum decrease in POP is 5.9$\times$ greater than Zeal because POP’s subproblems struggle to find alternative resources within their restricted subsets, whereas Zeal ’s subproblems retain access to the full set of resources. Teal exhibits a slight 1.3% increase in satisfied demand when the mean edge betweenness centrality is minimal. This is because the reduced edge selection simplifies allocation decisions.
@@
Figure 7(b) demonstrates that Zeal effectively manages temporal fluctuations, with changes in normalized satisfied demand limited to 1.3–3.5%. In contrast, Teal suffers a significant decrease of 12.8%, as it lacks prior exposure to the distribution of these fluctuating demands during training. Pinning and POP experience variations in normalized satisfied demand ranging from 0.6–3% and 1–1.6%, respectively. This is because randomness may occasionally favor splitting a problem into subproblems or focusing on top demands exclusively, while at other times, it does not.
Spatial redistribution.  To evaluate robustness against spatial redistribution, we reassign traffic demands among node pairs. Specifically, we adjust the top 10% of demands, which initially account for 88.4% of the total volume, so that they take up 80%, 60%, 40%, and 20% of the total volume instead. Figure 7(c) shows that Zeal consistently fulfills the highest demand across all spatial distributions. Pinning’s performance decreases by 8.3%, as its strategy of focusing only on top demands relies on the presence of a heavy-tailed demand distribution. Teal has a slight decrease up to 2.4% due to unseen demand distribution in training. This is a relatively smaller decrease than the temporal fluctuation, as Teal can still handle simple, evenly distributed traffic demands. POP has a slight increase of 1.4% as demands are more granular, where no demand requires a significant fraction of the total resources.
@@ 7.3 Micro-benchmarks
We conduct micro-benchmarks to evaluate the impact of Zeal’s design components, with all evaluations performed in maximizing total flow in traffic engineering.
Speedup when varying CPU cores.  Figure 9(a) evaluates the speedup of Zeal, Zeal, and Exact Sol with 1, 4, 16, and 64 CPU cores. Zeal achieves 61.7$\times$ speedup under 64 cores, where the slight gap occurs as we cannot exactly divide the total computation time evenly across 64 cores. Zeal also achieves an almost linear speedup when the number of CPU cores grows from 1 to 16. However, this linearity diminishes with 64 cores, where Zeal only achieves 18.2$\times$ speedup. This is because utilizing all 64 cores for parallelism incurs overhead from cache contention and stragglers, as discussed in Section 7.1. In contrast to the linear speedup of Zeal and Zeal, the speedup of Exact Sol is sublinear and marginal, achieving only 3.4$\times$ speeds with 64 CPU cores. This is because of commercial solvers’ sequential nature, e.g., the simplex method [45] in linear programming requires thousands to millions of small steps to reach the optimal solution.
Convergence rate.  Both Teal and Zeal employ ADMM for parallel optimization, but their approaches differ. Teal’s ADMM adapts its computations based on matrix multiplication to leverage GPU calculation capabilities. Figure 9(b) illustrates the convergence rates of Teal and Zeal under various initialization. Zeal takes 30 seconds to surpass 90% satisfied demand. However, when initialized with Teal’s coarse neural network solution, Zeal achieves this within 20 seconds. Conversely, with naive initialization, Zeal requires more than 60 seconds. Teal, benefiting from the massive parallelism of GPUs, reaches 90% satisfied demand within 10 seconds. However, its adapted ADMM limits its performance, capping its satisfied demand at 91% at 60 seconds compared to Zeal’s 92%.
Comparison to more optimization methods.  Figure 9(c) investigates more constrained optimization methods other than Zeal’s ADMM method in maximizing the total flow in traffic engineering. Section 7.1 has covered the computation time of the optimal Exact sol, the suboptimal POP-64, and the near-optimal Zeal. As Zeal replaces constraints with additional terms in the objective function, we compare similar methods, including the penalty method and the augmented Lagrangian method. The augmented Lagrangian method is the same as in Zeal’s ADMM but without decoupling and decomposition, and thus it is over 3$\times$ slower than Zeal in reaching a satisfied demand of over 90%. The penalty method, which gradually increases penalty parameters from 0 to infinity to enforce the converted constraints in the objective function, converges slowly and is over 30$\times$ slower than Zeal in reaching a satisfied demand of over 90%.
@@ 9 Conclusion
In this paper, we scale real-world resource allocation problems from a fresh perspective, observing that most of them exhibit an inherent separable structure. This key insight allows us to systematically decouple resource and demand constraints and decompose the original optimization problem into independent and parallel subproblems for allocating individual resources and demands. We implement this approach, Zeal, as a library to augment a widely-used, open-source solver (cvxpy), offering a push-button solution that accelerates large-scale resource allocation. Through extensive evaluation across diverse resource allocation tasks, we demonstrate that Zeal significantly improves scalability while closely approximating the quality of exact solutions.
Appendix A Mathematical Representation of Zeal

### v2

@@ 7.1.2 Cluster Scheduling
Max-min allocation.  Figure 5 compares DeDe against other approaches with the objective of max-min allocation. DeDe achieves a near-optimal max-min allocation of 0.97 within 5 seconds and 0.98 within 10 seconds, while DeDe reaches 0.98 within 2 seconds. Due to their sequential nature, Exact Sol and Gandiva encounter a slow computation time of 86 and 22 seconds. The greedy heuristic Gandiva even experiences a 0.56 low max-min allocation. POP-4 deliverers a comparable time-accuracy tradeoff as DeDe but is 3.4$\times$ slower than DeDe, which is simulated in the same way as POP did for a fair comparison. A faster POP variant, POP-16, exhibits a lower max-min allocation of 0.94, as it becomes more difficult for jobs to assess the ideal GPU/CPU among the limited selection in smaller subproblems.
@@
Proportional fairness.  Figure 6 compares DeDe against other approaches in maximizing proportional fairness. As the optimization problem becomes more difficult with the convex logarithm function in the objective, both DeDe and DeDe achieve a greater speedup than POP by thoroughly breaking down the problem into smaller and easier pieces. DeDe and DeDe reach a proportional fairness score of over 1,000 in 300 seconds and 100 seconds, respectively. In contrast, POP-4 takes 1,825 seconds to reach the same score, while POP-16 and POP-64 remain below 900 despite their shorter computation times. Exact Sol, though accurate, is prohibitively slow, requiring 12,432 seconds.
@@ 7.1.3 Load Balancing
Minimize shard movements.  Figure 8 compares DeDe with state-of-the-art approaches in minimizing shard movements. DeDe achieves an average of 24.1 shard movements in 15 seconds, slightly better than POP-4’s 25.4 shard movements within 16.9 seconds. In faster scenarios, DeDe achieves 35.1 shard movements in 2 seconds, while POP-4 achieves slightly fewer movements of 32.2 in 1.7 seconds. In a more fair comparison with POP, DeDe consistently costs 15–18% fewer shard movements than POP given a similar time limit. Exact Sol achieves the minimum shard movements of 13.2 but is significantly slower, taking an average of 1,572 seconds due to the NP-hardness of solving the mixed-integer linear programming problem inherent to load balancing. On the other hand, the greedy heuristic represents the opposite end of the time-accuracy trade-off, completing in just 25 milliseconds but requiring 50 shard movements.
@@ 7.2 Robustness
We evaluate the robustness of DeDe and other baselines against changes in problem granularity, temporal fluctuations, spatial redistribution, and link failures. In all these evaluations, DeDe is run for 30 seconds. To mitigate the impact of the changing resources and demands, we normalize the satisfied demand by the optimal satisfied demand from Exact Sol.
@@
As the mean edge betweenness centrality decreases, all approaches experience a decline in normalized satisfied demand, ranging from 0.8–2.1% for DeDe, 2.8–4.2% for Teal, 2.5–5.6% for Pinning, and 1–12.3% for POP. However, the maximum decrease in POP is 5.9$\times$ greater than DeDe because POP’s subproblems struggle to find alternative resources within their restricted subsets, whereas DeDe’s subproblems retain access to the full set of resources. Teal exhibits a slight 1.3% increase in satisfied demand when the mean edge betweenness centrality is minimal. This is because the reduced edge selection simplifies allocation decisions.
@@
Figure 7(b) demonstrates that DeDe effectively manages temporal fluctuations, with changes in normalized satisfied demand limited to 1.3–3.5%. In contrast, Teal suffers a significant decrease of 12.8%, as it lacks prior exposure to the distribution of these fluctuating demands during training. Pinning and POP experience variations in normalized satisfied demand ranging from 0.6–3% and 1–1.6%, respectively. This is because randomness may occasionally favor splitting a problem into subproblems or focusing on top demands exclusively, while at other times, it does not.
Spatial redistribution.  To evaluate robustness against spatial redistribution, we reassign traffic demands among node pairs. Specifically, we adjust the top 10% of demands, which initially account for 88.4% of the total volume, so that they take up 80%, 60%, 40%, and 20% of the total volume instead. Figure 7(c) shows that DeDe consistently fulfills the highest demand across all spatial distributions. Pinning’s performance decreases by 8.3%, as its strategy of focusing only on top demands relies on the presence of a heavy-tailed demand distribution. Teal has a slight decrease up to 2.4% due to unseen demand distribution in training. This is a relatively smaller decrease than the temporal fluctuation, as Teal can still handle simple, evenly distributed traffic demands. POP has a slight increase of 1.4% as demands are more granular, where no demand requires a significant fraction of the total resources.
@@ 7.3 Micro-benchmarks
We conduct micro-benchmarks to evaluate the impact of DeDe’s design components, with all evaluations performed in maximizing total flow in traffic engineering.
Speedup when varying CPU cores.  Figure 9(a) evaluates the speedup of DeDe, DeDe, and Exact Sol with 1, 4, 16, and 64 CPU cores. DeDe achieves 61.7$\times$ speedup under 64 cores, where the slight gap occurs as we cannot exactly divide the total computation time evenly across 64 cores. DeDe also achieves an almost linear speedup when the number of CPU cores grows from 1 to 16. However, this linearity diminishes with 64 cores, where DeDe only achieves 18.2$\times$ speedup. This is because utilizing all 64 cores for parallelism incurs overhead from cache contention and stragglers, as discussed in Section 7.1. In contrast to the linear speedup of DeDe and DeDe, the speedup of Exact Sol is sublinear and marginal, achieving only 3.4$\times$ speeds with 64 CPU cores. This is because of commercial solvers’ sequential nature, e.g., the simplex method [45] in linear programming requires thousands to millions of small steps to reach the optimal solution.
Convergence rate.  Both Teal and DeDe employ ADMM for parallel optimization, but their approaches differ. Teal’s ADMM adapts its computations based on matrix multiplication to leverage GPU calculation capabilities. Figure 9(b) illustrates the convergence rates of Teal and DeDe under various initialization. DeDe takes 30 seconds to surpass 90% satisfied demand. However, when initialized with Teal’s coarse neural network solution, DeDe achieves this within 20 seconds. Conversely, with naive initialization, DeDe requires more than 60 seconds. Teal, benefiting from the massive parallelism of GPUs, reaches 90% satisfied demand within 10 seconds. However, its adapted ADMM limits its performance, capping its satisfied demand at 91% at 60 seconds compared to DeDe’s 92%.
Comparison to more optimization methods.  Figure 9(c) investigates more constrained optimization methods other than DeDe’s ADMM method in maximizing the total flow in traffic engineering. Section 7.1 has covered the computation time of the optimal Exact sol, the suboptimal POP-64, and the near-optimal DeDe. As DeDe replaces constraints with additional terms in the objective function, we compare similar methods, including the penalty method and the augmented Lagrangian method. The augmented Lagrangian method is the same as in DeDe’s ADMM but without decoupling and decomposition, and thus it is over 3$\times$ slower than DeDe in reaching a satisfied demand of over 90%. The penalty method, which gradually increases penalty parameters from 0 to infinity to enforce the converted constraints in the objective function, converges slowly and is over 30$\times$ slower than DeDe in reaching a satisfied demand of over 90%.
@@ 9 Conclusion
In this paper, we scale real-world resource allocation problems from a fresh perspective, observing that most of them exhibit an inherent separable structure. This key insight allows us to systematically decouple resource and demand constraints and decompose the original optimization problem into independent and parallel subproblems for allocating individual resources and demands. We implement this approach, DeDe, as a library to augment a widely-used, open-source solver (cvxpy), offering a push-button solution that accelerates large-scale resource allocation. Through extensive evaluation across diverse resource allocation tasks, we demonstrate that DeDe significantly improves scalability while closely approximating the quality of exact solutions.
Appendix A Mathematical Representation of DeDe